OpenStudy (anonymous):

y=3x+7/3 how to solve this?

OpenStudy (anonymous):

anyone can help?

OpenStudy (anonymous):

hi rox13kh! :)

OpenStudy (anonymous):

hello what's the question ? ? are your suppose to solve for x ? ?

OpenStudy (anonymous):

yes but I think sth is missing

OpenStudy (anonymous):

*you try checking your source

OpenStudy (anonymous):

so there should be another data yes? for example y=5 and then I will be able to solve this

OpenStudy (anonymous):

so the question is solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aaa there you go... if y=5 then x=(5-7/3)/3 x=8/9

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

sure ^^ thx for the medal btw